OpenStudy (unklerhaukus):

Q: solve the DE \(y^{\prime\prime}+x(y^\prime)^2=0\) A: \(y=c\) or \(y=-\ln x+c\)

OpenStudy (unklerhaukus):

\[y^{\prime\prime}+x(y^\prime)^2=0\]\[\text{let } y^\prime=p\]\[y^{\prime\prime}=p^\prime\]\[p^\prime+xp^2=0\]\[p^\prime=-xp^2\]\[\frac{\text dp}{\text dx}=-x{p^2}\]\[\int\frac{{-\text dp}}{p^2}=\int x\text dx\]\[\frac{1}{p}=\frac{x^2}{2}+c\]\[\frac{\text dy}{\text dx}=\frac{2}{x^2+2c}\]\[{\text dy}=\frac{2}{x^2+2c}{\text dx}\]\[y=\int\frac{2}{x^2+2c}\text dx\]

OpenStudy (unklerhaukus):

im just a bit confuse with the constants, i think i need a negative to get logs when integrating

OpenStudy (experimentx):

$$ y = 2\tan^{-1}{\frac{x}{\sqrt{2C_1}}} + C_2$$

OpenStudy (unklerhaukus):

i am trying to get to the solution posted in the question box

OpenStudy (experimentx):

a second order DE must have two constants in it's general solution.

OpenStudy (unklerhaukus):

so the answers for the back of my book are incorrect @experimentX ?

OpenStudy (experimentx):

y = c is plainly incorrect.

OpenStudy (unklerhaukus):

0=0

OpenStudy (experimentx):

as for y=−lnx+c let's differentiate it dy/dx = -1/x <--- this is your DE which is of first order.

OpenStudy (experimentx):

y = c => dy/dx = 0

OpenStudy (experimentx):

wolf says http://www.wolframalpha.com/input/?i=y%27%27+%2B+x+%28y%27%29^2+%3D+0

OpenStudy (unklerhaukus):

yeah, and?

OpenStudy (experimentx):

I think that the given answer is quite not right.

OpenStudy (unklerhaukus):

well that is the answer i am trying to get to, what would you have the logarithmic form of the answer be?

OpenStudy (experimentx):

what do you mean by logarithmic form??

OpenStudy (unklerhaukus):

with logs rather than inverse trig identities

OpenStudy (experimentx):

Oh wait ..

OpenStudy (unklerhaukus):

the difference has something to do with the domain for x

OpenStudy (unklerhaukus):

@Zarkon

OpenStudy (experimentx):

for some strange reason wolf is saying \[ y = \frac {\sqrt 2} { \sqrt {c_1}}\tan^{-1}{\frac{x}{\sqrt{2C_1}}} + C_2 = y(x) = c_2-(\sqrt 2 \tanh^{-1}{x/(\sqrt 2 \sqrt{c_1}))}/\sqrt{c_1}\]

OpenStudy (phi):

Maybe I am missing something, but when I plug y=−lnx+c into the original differential equation, I don't get zero.

OpenStudy (unklerhaukus):

that is true @phi\[\frac 1{x^2}+x\times\left(-\frac1x\right)^2\neq0\] i guess it must be a mistake in the book , i have checked it multiple times,

OpenStudy (unklerhaukus):

but \[y=c\]\[0+x\times(0)^2=0\]

OpenStudy (experimentx):

well y = c is one solution but not complete solution.